# Neural Thermodynamic Integration: Free Energies from Energy-based Diffusion Models

### v2

@@ VI Conclusion
We present Neural TI: a generative machine-learning approach to perform thermodynamic integration in molecular systems. Our work shows that energy-based denoising diffusion models (DDMs) are particularly well suited to calculate the free-energy difference of turning on interactions in a many-body Hamiltonian. We find that by associating the latent space to the ideal-gas system, and further parametrizing the score as the derivative of an energy function, we can efficiently and accurately perform TI. Unlike conventional applications, our approach does not require reference Monte Carlo or molecular dynamics simulations at intermediate couplings between the end points. Instead, DDMs integrate the conformational ensemble along the finite-time interval of the diffusion process. Critically, we demonstrate for the first time the accurate estimation of free energies of coupling all interactions for up to 600 degrees of freedom.
@@
We demonstrate the applicability of Neural TI on Lennard-Jones fluids. We show our DDM model to transfer across densities around the gas–liquid transition of the system. The structural accuracy of the configurations is illustrated by the radial distribution functions. More importantly, we show that our TI calculations are accurate for varying numbers of particles: the excess chemical potentials and particle-number distributions result from weighted averages of TI-estimated canonical partition functions. We report free-energy differences of coupling Hamiltonians of up to $200\leavevmode\nobreak\ k_{\text{B}}T$ from a single reference simulation—the fully interacting Hamiltonian alone. We expect DDM-based TI to significantly improve alchemical transformations of molecular systems by coupling many more degrees of freedom than what was possible until now.
@@ Acknowledgements
We thank Fred Hamprecht and Daniel Nagel for fruitful discussions. BM acknowledges financial support by the Swiss National Science Foundation under grant number CR - SII5 - 193716 (RODEM). TB acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).
@@ Appendix A Experiments on a 1D Lennard-Jones system
Figure 7: Radial distribution functions as predicted by Monte Carlo simulations (gray) and a diffusion model (red) trained on densities $\rho\in\{0.5,0.7,0.9\}$. Note that the model reconstructs $g(r)$ on a wide range of densities starting from $\rho=0.4$, close to the gas phase to $\rho=1.0$ close to the solid phase. A Gaussian kernel with $\sigma=0.05$ was applied to all curves in this plot.
Figure 8: Distribution of the number of particles in the grand canonical ensemble at different chemical potentials as predicted by GCMC simulations (gray) and by thermodynamic integration with a trained diffusion model (red). The blue lines denote the canonical ensembles that the diffusion model was trained on. A Gaussian kernel with $\sigma=0.015$ was applied to all curves in this plot.
@@ Appendix B Experimental details
Force fields
@@
For the $1$-dimensional box (Appendix A), we use a SchNet-like [28] architecture. To make the architecture time-dependent, we predict the parameters of the RBF kernels from the diffusion time using a small MLP. The network consists of 3 layers each of which performs a message passing step and an atom-wise update. The final readout of the energy is the sum of the features over all nodes and channels. The number of channels is 32 and the MLPs in the RBF-parameter prediction and in the atom-wise update both have two hidden layers with 64 neurons. Since the number of particles is reasonably low ($\leq 100$), we do not use a cutoff radius, and perform message passing on the fully connected graph between the nodes.

### v3

@@ Conclusion
We present Neural TI: a generative machine-learning approach to perform thermodynamic integration in molecular systems. Our work shows that energy-based denoising diffusion models (DDMs) are particularly well suited to calculate the free-energy difference of turning on interactions in a many-body Hamiltonian. We find that by associating the latent space to the ideal-gas system, and further parametrizing the score as the derivative of an energy function, we can efficiently and accurately perform TI. Unlike conventional applications, our approach does not require reference Monte Carlo or molecular dynamics simulations at intermediate couplings between the end points. Instead, DDMs integrate the conformational ensemble along the finite-time interval of the diffusion process. We demonstrate the accurate estimation of the free-energy difference of coupling the LJ interactions from a box of ideal-gas particles.
@@
We demonstrate the applicability of Neural TI on Lennard-Jones fluids. We show our DDM model to transfer across densities around the gas–liquid transition of the system. The structural accuracy of the configurations is illustrated by the radial distribution functions. More importantly, we show that our TI calculations are accurate for varying numbers of particles: the excess chemical potentials and particle-number distributions result from weighted averages of TI-estimated canonical partition functions. We report free-energy differences of coupling Hamiltonians of up to $200\leavevmode\nobreak\ k_{\text{B}}T$ from a single reference simulation—the fully interacting Hamiltonian alone.
@@ Acknowledgements
We thank Fred Hamprecht and Daniel Nagel for fruitful discussions. We also thank Chin-Wei Huang for bringing the work of Masrani et al. [22] to the authors’ attention. BM acknowledges financial support by the Swiss National Science Foundation under grant number CR - SII5 - 193716 (RODEM). TB acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).
Data availability
@@ Appendix A Experiments on a 1D Lennard-Jones system
Figure 7: Radial distribution functions as predicted by Monte Carlo simulations (gray) and a diffusion model (red) trained on densities $\rho\in\{0.5,0.7,0.9\}$. Note that the model reconstructs $g(r)$ on a wide range of densities starting from $\rho=0.4$, close to the gas phase to $\rho=1.0$ close to the solid phase.
Figure 8: Distribution of the number of particles in the grand canonical ensemble at different chemical potentials as predicted by GCMC simulations (gray) and by thermodynamic integration with a trained diffusion model (red). The blue lines denote the canonical ensembles that the diffusion model was trained on.
@@ Appendix B Experimental details
TI in Figure 6
We run simulations at $\lambda\in\{0.00,0.01,0.02,...,1.00\}$ at particle counts $10,20,...,200$ with the intermediate potential
$U_{\lambda}^{\text{LJ}}(\mathbf{q})=\sum_{i\neq j}4(1-\lambda)\varepsilon\left[\left(\frac{\sigma^{2}}{\lambda\sigma^{2}+r_{ij}^{2}}\right)^{6}-\left(\frac{\sigma^{2}}{\lambda\sigma^{2}+r_{ij}^{2}}\right)^{3}\right].$
We collect $12000$ samples from which discard the first $2000$ as the burn-in phase.
Implementation details
We minimize a noise prediction objective $\mathbb{E}_{t\sim\mathcal{U}(0,1)}\mathbb{E}_{x_{0}\sim\rho_{0}}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}||\epsilon_{\theta}(x_{t},t)-\epsilon||^{2}$. We use the Adam optimizer with its learning rate starting from $10^{-3}$ and exponentially decayed to $10^{-5}$ over 100,000 training steps. The noise schedule of the DDM is chosen to be $\sigma_{t}=\sigma_{min}^{1-t}\sigma_{max}^{t}$ with $\sigma_{min}=10^{-3}$ and $\sigma_{max}=0.5$. The value of $\sigma_{min}$ was chosen to be much smaller than the length scale of the interaction potential, whereas $\sigma_{max}$ was set to be comparable to the size of the simulation box. At sampling time, potential issues of “lagging behind” can appear, but are quite distinct from, say, a system driven out of equilibrium by means of a time-dependent Hamiltonian. To mitigate this, we integrate the reverse dynamics with $1000$ time steps.
@@
For the $1$-dimensional box we use a SchNet-like [32] architecture. To make the architecture time-dependent, we predict the parameters of the RBF kernels from the diffusion time using a small MLP. The network consists of 3 layers each of which performs a message passing step and an atom-wise update. The final readout of the energy is the sum of the features over all nodes and channels. The number of channels is 32 and the MLPs in the RBF-parameter prediction and in the atom-wise update both have two hidden layers with 64 neurons. Since the number of particles is reasonably low ($\leq 100$), we do not use a cutoff radius, and perform message passing on the fully connected graph between the nodes.